# Fast Sampling via Discrete Non-Markov Diffusion Models

### v2

@@ Abstract
Discrete diffusion models have emerged as powerful tools for high-quality data generation. Despite their success in discrete spaces, such as text generation tasks, the acceleration of discrete diffusion models remains under explored. In this paper, we propose a discrete non-Markov diffusion model, which admits an accelerated reverse sampling for discrete data generation. Our method significantly reduces the number of function evaluations (i.e., calls to the neural network), making the sampling process much faster. Furthermore, we study the transition from finite to infinite step sampling, offering new insights into bridging the gap between discrete and continuous-time processes for discrete diffusion models. Extensive experiments on natural language generation and machine translation tasks demonstrate the superior performance of our method in terms of both generation speed and sample quality compared to existing methods for discrete diffusion models.
@@ 1 Introduction
  • We propose discrete non-Markov diffusion models (DNDM), which preserve two important properties of the original discrete Markov diffusion model. Firstly, for any diffusion trajectory $\{\mathbf{x}_{t}\}$ starting from real data $\mathbf{x}_{0}$, DNDM provably preserves the marginal distribution $q(\mathbf{x}_{t})$ and the conditional distribution $q(\mathbf{x}_{0}|\mathbf{x}_{t})$. Secondly, the proposed DNDM model naturally induces a set of latent variables $\mathcal{T}$, termed as the transition time set. A predetermined transition time set allows us to develop a training-free sampling algorithm that can accelerate a large family of discrete diffusion models, including the two most widely used discrete diffusion models: multinomial diffusion (Hoogeboom et al., 2021b) and absorbing diffusions (Austin et al., 2021). DDIM is non-Markov and admits a de-randomized, faster sampling algorithm compared to DDPM. In this context, DNDM can be considered as the discrete counterpart of DDIM (See Table 1).
  • Based on the DNDM, we design an accelerated sampling algorithm that reduces the required number of neural network function evaluations. In a standard $T$ time-step discrete diffusion process, while D3PM, including Multinomial (Ho et al., 2020) and absorbing state discrete sampling (Austin et al., 2021), requires evaluating the neural network function $T$ times, our approach only requires $|{\mathcal{T}}|$ function evaluations, where $|{\mathcal{T}}|$ is the cardinality of the transition set ${\mathcal{T}}$. Moreover, $|{\mathcal{T}}|$ is provably less than $T$ and approaches $O(1)$ as $T$ goes to infinity. We provide both theoretical analysis and empirical experiments showing that the improvement in the number of function evaluations (NFE) is significant. Notably, our algorithm is about 3×3$\times$3 × faster than baselines for $T=50$ and about 30×30$\times$30 × faster for $T=1000$ while preserving the sample quality.
@@ 3.3 Continous-time (Infinite Step) Reverse Sampling
for all $n\in[N]$. The detailed algorithm named DNDM-C is shown in Algorithm 4 in Appendix G.1.
@@ 4.1 Conditional Text Generation
We first evaluate the ability of DNDM on conditional text generation by machine translation tasks. Following (Zheng et al., 2023), we process the raw text with the Byte Pair Encoder (BPE) (Sennrich et al., 2016) to construct the vocabulary, which consists of the words and subwords of both the source and the target languages. We conduct our experiments using the FairSeq toolkit (Ott et al., 2019), which employs a model consisting of an encoder and a decoder. The encoder takes the source text as input, while the decoder generates the target text.
@@
Baselines. The main baselines we are comparing with are RDM and RDM-$k$ from Zheng et al. (2023). Here, we use RDM and RDM-$k$ to denote the sampling method proposed in their paper with and without the usage of top-$k$ selection for the token generation technique (see Appendix E for more details), respectively. RDM and RDM-$k$ are applied to two previously proposed state-of-the-art discrete diffusion models: Multinomial Diffusion (Hoogeboom et al., 2021b) and Absorbing Diffusion (Austin et al., 2021).
@@
Scaling Law in Sampling Speed. For illustrative purposes, we use the example of $\mathtt{IWSLT14}$ to visualize how the sample quality scales regarding sampling speed for different methods. In Figure 1, we observe the trend of the BLEU score in relation to computational time. Each line in the legend represents a different sampling algorithm, and a steeper slope indicates a larger marginal gain when sampling for longer periods. Figure 1 demonstrates that our algorithm displays nearly linear growth in BLEU score over the log of time, which is remarkable in contrast with the flat curve of the baseline. Particularly, for multinomial diffusion, the BLEU score increases by 1 in less than 60 seconds of additional sampling time. For absorbing diffusion, DNDM outperforms RDM before RDM samples 50 steps. In Tables 6 and 7 in Appendix D, we further use the average number of function evaluations (NFE) to measure the improved speed within the specified number of sampling steps. Additionally, in Figure 2, we visualize how the BLEU score and the generated text change throughout the sampling process.
@@ 4.2 Unconditional Text Generation
In unconditional text generation, we focus on language modeling tasks, where the goal is to generate language data similar to the provided training dataset. In this task, no input text is given during sampling, and the neural network directly learns $q(\mathbf{x}_{0}|\mathbf{x}_{t})$.
@@ 5 Conclusion and Future Work
This paper presents a novel discrete non-Markov diffusion model (DNDM) accompanied by an accelerated sampling algorithm designed to boost sampling speed in a discrete-state space. Our discrete diffusion model incorporates "transition time set" latent variables, establishing itself as an efficacious diffusion and data generation method. Thanks to our acceleration technique, we significantly decrease the number of neural network function evaluations without sacrificing sample quality. We also introduce an infinite-step sampling algorithm, DNDM-C, which provides new insights into bridging the gap between discrete and continuous-time processes for discrete diffusion models. While this study focuses on text generation using non-autoregressive models, a promising direction for future exploration is applying our method to other tasks, such as audio and image generation, as well as other architectures, such as the GPT model (an autoregressive model).
@@ Appendix C Choice of the Transition Time
Additionally, the ablation study on a reasonable range of different Beta distributions with 50 and 1000 sampling steps are shown in Tables 9 and 8, where the BLEU scores and NFE values on the test set of one of the three machine translation datasets, $\mathtt{WMT16}$, are shown for demonstration. The range of Beta distributions covers our chosen Beta schedules based on validation sets and a variety of basic Beta distribution shapes. These results show that the different Beta distributions influence the performance, but most of these choices of parameters still achieve results close to the optimal. Since the Beta distributions of the reported results in Tables 2 and 3 are selected using the validation set, they do not always have the highest scores on the test set, but their scores still at least belong to the top tiers according to these tables.
@@
In practical applications of our model, we employ a beta distribution for $\mathcal{D}_{\tau}$, which typically exhibits a right-heavy tail. Therefore $C_{T,N,\mathcal{D}_{\tau}}$ tends to be larger than that in the worst-case scenario. In Tables 6 and 7, we list the average NFE for each experiment we run in §4. These results demonstrate a significant reduction in NFE compared to the original counts: for $T=25$, the NFE is only about half of the original count; for $T=50$, it is approximately one-third; and for $T=1000$, it reduces to less than one-twentieth of the original count.
@@ F.1 Conditional Text Generation
We conduct a performance comparison based on varying configurations of the Beta and Alpha distributions. The results of these comparisons are presented in Tables 9 and 8. Furthermore, to evaluate the efficacy of discrete versus continuous step schemes, we also conduct an ablation study under the same set of parameters $(100,4)$ in Table 10.
@@ G.1 Continuous Training
Comparison with ARDM (Hoogeboom et al., 2021a). Autoregressive Diffusion Model (ARDM) (Hoogeboom et al., 2021a) is a discrete diffusion model built upon the autoregressive nature of data. ARDM is shown to be equivalent to a continuous-time absorbing diffusion model and thus provides a unique perspective for discrete diffusion. For continuous-time ($T=\infty$) reverse sampling, both ARDM and our method achieve $N$ NFEs. Compared with ARDM, our method provides a unified framework including both absorbing and multinomial diffusions, applicable to both finite time and continuous time diffusions. For infinite timesteps, Hoogeboom et al. (2021a) also proposed an advanced parallelizing technique that can reduce NFE according to the log-likelihood, which we have not considered in DNDM-C.
Algorithm 4 Sampling from DNDM-C
1:  for $n=1\ldots N$ do
2:     Initiate each token $\mathbf{x}_{T,n}\sim\mathbf{q}_{\mathrm{noise}}$
4:  end for
5:  for $k=N\ldots 1$ do
7:     for $n=1\ldots N$ do
8:        Update $\mathbf{x}_{\tau_{n_{k-1}},n}$ based on condition of $\tau_{n}$
9:     end for
10:  end for
11:  Return $\mathbf{x}_{0,1:N}$

### v3

@@ Abstract
Discrete diffusion models have emerged as powerful tools for high-quality data generation. Despite their success in discrete spaces, such as text generation tasks, the acceleration of discrete diffusion models remains under-explored. In this paper, we propose discrete non-Markov diffusion models (DNDM), which naturally induce the predetermined transition time set. This enables a training-free sampling algorithm that significantly reduces the number of function evaluations (i.e., calls to the neural network), making the sampling process much faster. Furthermore, we study the transition from finite to infinite step sampling, offering new insights into bridging the gap between discrete and continuous-time processes for discrete diffusion models. Extensive experiments on natural language generation and machine translation tasks demonstrate the superior performance of our method in terms of both generation speed and sample quality compared to existing methods for discrete diffusion models. Codes are available at https://github.com/uclaml/DNDM.
@@ 1 Introduction
  • We propose discrete non-Markov diffusion models (DNDM), which naturally induces a set of latent variables $\mathcal{T}$, termed as the transition time set. This key feature enables us to develop a training-free sampling algorithm that can accelerate a large family of discrete diffusion models. Importantly, DNDM preserves the essential properties of the original discrete diffusion model: for any diffusion trajectory $\{\mathbf{x}_{t}\}$ starting from real data $\mathbf{x}_{0}$, it provably maintains both the marginal distribution $q(\mathbf{x}_{t})$ and the conditional distribution $q(\mathbf{x}_{0}|\mathbf{x}_{t})$. Our method can accelerate the two most widely used discrete diffusion models: multinomial diffusion (Hoogeboom et al., 2021b) and absorbing diffusions (Austin et al., 2021). Similar to how DDIM introduces a de-randomized, faster sampling algorithm compared to DDPM in continuous space, DNDM achieves acceleration through a predetermined transition time set in discrete space (See Table 1).
  • Based on the predetermined transition time set ${\mathcal{T}}$ in DNDM, we design an accelerated sampling algorithm that reduces the required number of neural network function evaluations. In a standard $T$ time-step discrete diffusion process, while D3PM, including Multinomial (Ho et al., 2020) and absorbing state discrete sampling (Austin et al., 2021), requires evaluating the neural network function $T$ times, our approach only requires $|{\mathcal{T}}|$ function evaluations, where $|{\mathcal{T}}|$ is the cardinality of the transition set ${\mathcal{T}}$. Moreover, $|{\mathcal{T}}|$ is provably less than $T$ and approaches $O(1)$ as $T$ goes to infinity. We provide both theoretical analysis and empirical experiments showing that the improvement in the number of function evaluations (NFE) is significant. Notably, our algorithm is about 3×3$\times$3 × faster than baselines for $T=50$ and about 30×30$\times$30 × faster for $T=1000$ while preserving the sample quality.
@@ 3.1 Forward and Reverse Process
Remark 3.5.
While Song et al. (2020a) proposed a non-Markov multinomial diffusion model in Appendix A, DDIM and DNDM are fundamentally different models when specialized to multinomial diffusion. DDIM’s discrete process remains stochastic at every step, even with deterministic noise scheduling. In contrast, DNDM achieves full de-randomization by pre-determined transition time $\tau$ (Equation 8 in our paper). By sampling these transition times upfront, DNDM establishes a predetermined transition time set that guides the sampling process, enabling deterministic evolution and faster sampling speed even under the same number of sampling steps, which is not reported under DDIM framework. For detailed technical comparison, see Appendix B.1.
@@ 3.3 Continous-time (Infinite Step) Reverse Sampling
Algorithm 2 Sampling from DNDM-C
1:  for $n=1\ldots N$ do
2:     Initiate each token $\mathbf{x}_{T,n}\sim\mathbf{q}_{\mathrm{noise}}$
4:  end for
5:  for $k=N\ldots 1$ do
7:     for $n=1\ldots N$ do
8:        Update $\mathbf{x}_{\tau_{n_{k-1}},n}$ based on condition of $\tau_{n}$
9:     end for
10:  end for
11:  Return $\mathbf{x}_{0,1:N}$
@@
for all $n\in[N]$. The detailed algorithm named DNDM-C is shown in Algorithm 2.
Remark 3.7.
Autoregressive Diffusion Model (ARDM) (Hoogeboom et al., 2021a) is a discrete diffusion model built upon the autoregressive nature of data. ARDM is shown to be equivalent to a continuous-time absorbing diffusion model and thus provides a unique perspective for discrete diffusion. For continuous-time ($T=\infty$) reverse sampling, both ARDM and our method achieve $N$ NFEs. Unlike ARDM which is limited to absorbing-state transitions, our method provides a unified framework including both absorbing and multinomial diffusions, applicable to both finite time and continuous time diffusions. For infinite timesteps, Hoogeboom et al. (2021a) also proposed an advanced parallelizing technique that can reduce NFE according to the log-likelihood, which we have not considered in DNDM-C.
@@ 4.1 Conditional Text Generation
We evaluate DNDM’s effectiveness on conditional text generation through machine translation tasks. Following Zheng et al. (2023), we use Byte Pair Encoding (BPE) (Sennrich et al., 2016) to create a shared vocabulary of words and subwords from both source and target languages. We implement our experiments using FairSeq (Ott et al., 2019), which employs an encoder-decoder architecture. The model uses bi-directional self-attention blocks without causal masking, allowing tokens to attend to both past and future positions during training and inference. The encoder processes the source text, while the decoder generates the target translation.
@@
Baselines. The main baselines we are comparing with are RDM and RDM-$k$ from Zheng et al. (2023). Here, we use RDM-$k$ and RDM to denote the sampling method proposed in their paper with and without the usage of top-$k$ selection for the token generation technique (see Appendix E for more details), respectively. RDM and RDM-$k$ are applied to two previously proposed state-of-the-art discrete diffusion models: Multinomial Diffusion (Hoogeboom et al., 2021b) and Absorbing Diffusion (Austin et al., 2021).
@@
Scaling Law in Sampling Speed. For illustrative purposes, we use the example of $\mathtt{IWSLT14}$ to visualize how the sample quality scales regarding sampling speed for different methods. In Figure 1, we observe the trend of the BLEU score in relation to computational time. Each line in the legend represents a different sampling algorithm, and a steeper slope indicates a larger marginal gain when sampling for longer periods. Figure 1 demonstrates that our algorithm displays nearly linear growth in BLEU score over the log of time, which is remarkable in contrast with the flat curve of the baseline. Particularly, for multinomial diffusion, the BLEU score increases by 1 in less than 60 seconds of additional sampling time. For absorbing diffusion, DNDM outperforms RDM before RDM samples 50 steps. In Tables 7 and 8 in Appendix D, we further use the average number of function evaluations (NFE) to measure the improved speed within the specified number of sampling steps. Additionally, in Figure 2, we visualize how the BLEU score and the generated text change throughout the sampling process.
@@ 4.2 Unconditional Text Generation
For unconditional text generation, we evaluate our approach on language modeling tasks, where the model learns to generate text that matches the statistical patterns of the training data. Unlike conditional generation, this task involves directly learning $q(\mathbf{x}_{0}|\mathbf{x}_{t})$ without conditioning on any input text. We conduct experiments on the text8 and enwik8 datasets using a decoder-only architecture similar to GPT models. Since unconditional generation does not require encoding input sequences, we employ a 12-layer Transformer decoder without an encoder component.
@@ 5 Conclusion and Future Work
This paper presents a novel discrete non-Markov diffusion model (DNDM) accompanied by an accelerated sampling algorithm designed to boost sampling speed in a discrete-state space. Our discrete diffusion model incorporates "transition time set" latent variables, establishing itself as an efficacious diffusion and data generation method. Thanks to our acceleration technique, we significantly decrease the number of neural network function evaluations without sacrificing sample quality. We also introduce an infinite-step sampling algorithm, DNDM-C, which provides new insights into bridging the gap between discrete and continuous-time processes for discrete diffusion models. While this study focuses on text generation using non-autoregressive models, a promising direction for future exploration is applying our method to other tasks, such as audio and image generation.
Acknowledgement
We thank the anonymous reviewers and area chair for their helpful comments. ZC, HY, YL, YK, JZ, and QG are supported in part by the National Science Foundation CAREER Award 1906169, IIS-2008981, and the Sloan Research Fellowship. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ B.1 Comparison between D3PM and DNDM
Comparison between D3PM-Absorb and DNDM. Recall the forward processes of D3PM and DNDM as follows:
[…]
$\mathrm{DNDM:}\quad\mathbf{x}_{t}$ $=b_{t}\mathbf{x}_{t-1}+(1-b_{t})\mathbf{w},\quad\forall t=1\dots T.$
For absorbing diffusion where $\mathbf{w}=\mathrm{[Mask]}$, DNDM’s forward process becomes equivalent to D3PM since $\mathbf{w}_{t}=\mathbf{w}=\mathrm{[Mask]}$ in this special case. However, for multinomial diffusion or other diffusion processes where $\mathbf{w}_{t}\neq\mathbf{w}$, these two processes exhibit different behaviors. In addition, even for absorbing diffusion, our proposed reverse sampling algorithm for DNDM is still different from that for D3PM.
To elucidate the key differences between the sampling algorithm in DNDM and that in D3PM for absorbing diffusion, let’s directly compare the algorithms:
  • For the D3PM-Absorb algorithm: We begin with an all $\mathrm{[Mask]}$ sequence. At each time step $t$, we sample $\mathbf{x}_{0}\sim p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})$. If $\mathbf{x}_{t}=\mathrm{[Mask]}$, $\mathbf{x}_{t-1}$ transitions to $\mathrm{[Mask]}$ with probability $(1-\alpha_{t-1})/(1-\alpha_{t})$ and to $\mathbf{x}_{0}$ with probability $(\alpha_{t-1}-\alpha_{t})/(1-\alpha_{t})$. If $x_{t}\not=\mathrm{[Mask]}$, it remains unchanged.
  • For the DNDM-Absorb algorithm: We also start with an all $\mathrm{[Mask]}$ sequence, but crucially, we first determine the transition time set. During sampling, if $\mathbf{x}_{t}=\mathrm{[Mask]}$, the transition probabilities for $\mathbf{x}_{t-1}$ are identical to D3PM. However, we only sample $\mathbf{x}_{0}\sim p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{t})$ when at least one token needs to change, as determined by our pre-computed transition set. This selective sampling is the key to our algorithm’s efficiency.
Therefore, you can see that DNDM will skip many steps during the sampling process to avoid function evaluation and save computational cost. Even though the forward process of DNDM is the same as that of D3PM for absorbing diffusion, our DNDM approach introduces an algorithm design in the sampling process by pre-computing the transition time set and selectively applying function evaluations. This distinguishes DNDM from D3PM algorithm, offering a more computationally efficient approach to inference in discrete diffusion.
Comparison between DDIM and DNDM for Multinomial Diffusion. While there are similarities between DNDM and DDIM (Appendix A), they are fundamentally different models, and DNDM is not a special case of DDIM. DNDM introduces a novel framework specifically designed for discrete spaces, while DDIM was originally developed for continuous diffusion models. The key differences for multinomial diffusion are as follows.
  • DDIM: Following Song et al. (2020a) (eq. 19 in Appendix A), $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\text{Cat}(\sigma_{t}\mathbf{x}_{t}+(\alpha_{t-1}-\sigma_{t}\alpha_{t})\mathbf{x}_{0}+((1-\alpha_{t-1})-(1-\alpha_{t})\sigma_{t})\mathbf{1}_{K})$. Even with $\sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}$, the process remains stochastic: $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\text{Cat}(\sigma_{t}\mathbf{x}_{t}+(1-\sigma_{t})\mathbf{x}_{0})$. This means at every step, there’s a probability of choosing $\mathbf{x}_{0}$, regardless of whether it has transitioned to $\mathbf{x}_{0}$ or not. Unlike Absorbing discrete diffusion, no $\mathrm{[Mask]}$ exists in multinomial diffusion. Therefore, DDIM cannot distinguish whether $\mathbf{x}_{t}$ already equals $\mathbf{x}_{0}$ or not. In particular, although the sampling process becomes less stochastic in the DDIM setting, it will still be predicted $\mathbf{x}_{0}$ with high probability $1-\sigma_{t}=\frac{\alpha_{t-1}-\alpha_{t}}{1-\alpha_{t}}$.
  • DNDM: Achieves full de-randomization using transition time $\tau$, where:
    $\mathbf{x}_{t-1}=\operatorname{\mathds{1}}(\tau=t)\mathbf{x}_{0}+\operatorname{\mathds{1}}(\tau\not=t)\mathbf{x}_{t},\quad\text{with }P(\tau=t)=\alpha_{t-1}-\alpha_{t}.$ (14)
    This crucial difference allows DNDM to achieve full de-randomization once $\tau$ is sampled, leading to a deterministic evolution that DDIM cannot achieve.
While DNDM and DDIM are both non-Markov models for multinomial diffusion, their fundamental approaches to and achievements in de-randomization differ significantly in discrete spaces.
@@ Appendix C Choice of the Transition Time
Additionally, the ablation study on a reasonable range of different Beta distributions with 50 and 1000 sampling steps are shown in Tables 10 and 9, where the BLEU scores and NFE values on the test set of one of the three machine translation datasets, $\mathtt{WMT16}$, are shown for demonstration. The range of Beta distributions covers our chosen Beta schedules based on validation sets and a variety of basic Beta distribution shapes. These results show that the different Beta distributions influence the performance, but most of these choices of parameters still achieve results close to the optimal. Since the Beta distributions of the reported results in Tables 2 and 3 are selected using the validation set, they do not always have the highest scores on the test set, but their scores still at least belong to the top tiers according to these tables.
@@
Impact of Transition Order. We further evaluate the impact of transition order. Building upon the results in Table 3, we investigate how the model performance will change if the transition time is influenced by the position of the tokens: from left to right and from right to left. In the left-to-right approach, tokens positioned on the left are transitioned to $\mathbf{x}_{0}$ earlier, and vice versa for the right-to-left approach. Our experiments show that the left-to-right approach consistently outperforms the right-to-left approach across all datasets and step counts, as demonstrated in Table 6.
Table 6: Comparison of left-to-right and right-to-left transition approaches across different datasets and step counts.
Steps Direction IWSLT14 WMT14 WMT16
25 Left-to-right 31.08 24.41 31.67
Right-to-left 30.54 23.33 31.33
50 Left-to-right 32.87 26.46 33.37
Right-to-left 32.47 25.18 32.78
1000 Left-to-right 34.45 27.93 34.43
Right-to-left 34.04 27.02 34.15
This result suggests that the order of token transitions significantly influences the model’s performance, with earlier transitions of left-side tokens leading to better generation quality.
@@
In practical applications of our model, we employ a beta distribution for $\mathcal{D}_{\tau}$, which typically exhibits a right-heavy tail. Therefore $C_{T,N,\mathcal{D}_{\tau}}$ tends to be larger than that in the worst-case scenario. In Tables 7 and 8, we list the average NFE for each experiment we run in §4. These results demonstrate a significant reduction in NFE compared to the original counts: for $T=25$, the NFE is only about half of the original count; for $T=50$, it is approximately one-third; and for $T=1000$, it reduces to less than one-twentieth of the original count.
@@ F.1 Conditional Text Generation
We conduct a performance comparison based on varying configurations of the Beta and Alpha distributions. The results of these comparisons are presented in Tables 10 and 9. Furthermore, to evaluate the efficacy of discrete versus continuous step schemes, we also conduct an ablation study under the same set of parameters $(100,4)$ in Table 11.